a ladder is set to rest against a very slippery wall at an angle of 55 °. How large must the coefficient of friction between the ladder and the floor be in order for the ladder to remain upright?


1
Expert's answer
2021-01-14T12:51:06-0500

Let’s consider the forces acting on the ladder: the weight of the ladder "mg" directed downward, the normal force "F_W" directed from the wall, the normal force "F_N" directed from the floor and the force of friction "F_{fr}" directed toward the wall. In order to ladder remains upright and not slipping, the sum of these forces must be equal to zero. Applying Newton’s laws we get:


"F_{fr}=F_W,""F_N=mg."


The sum of moments of forces around the pivot point at the bottom of the ladder equals:


"\sum \tau=0,""\tau_{ladder}-\tau_{wall}=0,""F_WLsin\theta-\dfrac{L}{2}mgcos\theta=0,""\mu_smgsin\theta-\dfrac{1}{2}mgcos\theta=0,""\mu_s=\dfrac{1}{2}cot\theta=\dfrac{1}{2}cot55^{\circ}=0.35"

Answer:

"\mu_s=0.35"